# On the analysis of a higher-order Lotka-Volterra model: an application of $\mathcal{S}$-tensors and the polynomial complementarity problem

### v1

@@ Abstract
It is known that the effect of species’ density on species’ growth is non-additive in real ecological systems. This challenges the conventional Lotka-Volterra model, where the interactions are always pairwise and their effects are additive. To address this challenge, we introduce HOIs (Higher-Order Interactions) which are able to capture, for example, the indirect effect of one species on a second one correlating to a third species. Towards this end, we propose a general higher-order Lotka-Volterra model. We provide an existence result of a positive equilibrium for a non-homogeneous polynomial equation system with the help of $\mathcal{S}$-tensors. Afterward, by utilizing the latter result, as well as the theory of monotone systems and results from the polynomial complementarity problem, we provide comprehensive results regarding the existence, uniqueness, and stability of the corresponding equilibrium. These results can be regarded as natural extensions of many analogous ones for the classical Lotka-Volterra model, especially in the case of full cooperation, competition among two factions, and pure competition. Finally, illustrative numerical examples are provided to highlight our contributions.
Index Terms:
@@ I Introduction
Since the Lotka-Volterra model is a polynomial system, computing a positive equilibrium is equivalent to solving a system of polynomial equations. With the development of the tensor algebra, a polynomial equation system can be efficiently captured by the tensor-vector product [14, 15, 16]. Researchers have mainly focused on homogeneous polynomial systems and have further shown that such systems may possess a unique positive solution for some particular classes of tensors and under some appropriate conditions. For example, [14] provides a result of the existence and uniqueness of a positive solution for $\mathcal{M}$-tensors, while [15] extends the results to $\mathcal{H}^{+}$-tensors, and [16] to $\mathcal{S}$-tensors, see section II. However, most of these results are restricted to a homogeneous polynomial case and are not directly applicable to the Lotka-Volterra model.
@@
Notation: Throughout this paper, $\mathbb{R}$ denotes the set of real numbers, $\mathbb{R}_{+}$ is the set of nonnegative real numbers, and $\mathbb{R}_{++}$ the set of positive real numbers. For the similar notation $\mathbb{R}^{n}$ (+$\mathbb{R}^{n}_{+}$ or ++$\mathbb{R}^{n}_{++}$), the superscript stands for the dimension of the space. Whenever $a,b\in\mathbb{R}^{n}$, we use the notation $a\geq(\leq)b$ to denote that $a_{i}\geq(\leq)b_{i}$, for all $i=1,\ldots,n$; $a>(<)b$ to denote that $a_{i}>(<)b_{i}$, for all $i=1,\ldots,n$. This element-wise comparison is also used for matrices and tensors. For simplicity, the equilibrium point denotes both the point itself and the vector constructed from the point. For a vector $X=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}$, the vector’s $L_{\infty}$ norm is defined as $\|X\|_{\infty}=\max_{i}\left|x_{i}\right|$.
II Tensor-based approach
In this section, we briefly introduce the concepts of tensors and polynomial complementarity problems that are useful later in the paper. We also propose some novel results, which are extensions of the known results in tensor algebra. We will then show that all these results are very useful tools to analyze the higher-order Lotka-Volterra system.
II-A Tensor
A tensor $T\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{k}}$ is a multidimensional array, where the order is the number of its dimensions $k$ and each dimension $n_{i}$, $i=1,\cdots,k$ is a mode of the tensor. A tensor is cubical if every mode has the same size, that is $T\in\mathbb{R}^{n\times n\times\cdots\times n}$. From now on we denote a $k$-th order $n$ dimensions cubical tensor as $T\in\mathbb{R}^{n\times n\times\cdots\times n}=\mathbb{R}^{[k,n]}$. Throughout this paper, a ‘tensor’ always refers to a cubical tensor. A tensor $T$ is called supersymmetric if $T_{j_{1}j_{2}\ldots j_{k}}$ is invariant under any permutation of the indices. The identity tensor $\mathcal{I}=\left(\delta_{i_{1}\cdots i_{m}}\right)$ is defined by
$\delta_{i_{1}\cdots i_{m}}=\begin{cases}1&\text{ if }i_{1}=i_{2}=\cdots=i_{m}\\ 0&\text{ otherwise }\end{cases}.$
@@ II-A Tensor
II-B $\mathcal{M}$-tensors and $\mathcal{H}^{+}$-tensors
Here, we briefly recall the concepts of $\mathcal{M}$-tensor and $\mathcal{H}^{+}$-tensor introduced in [15, 26, 27]. A tensor $A\in\mathbb{R}^{[m,n]}$ is an $\mathcal{M}$-tensor if it can be represented as $A=s\mathcal{I}-B$, where $\mathcal{I}\in\mathbb{R}^{[m,n]}$ is the identity tensor, $B\in\mathbb{R}^{[m,n]}$ is a nonnegative tensor (i.e., each entry of $B$ is nonnegative), and $s\geq\rho(B)$. Furthermore, $A$ is called a nonsingular $\mathcal{M}$-tensor if $s>\rho(B)$. A tensor $A\in\mathbb{R}^{[m,n]}$ is a Metzler tensor if it can be represented as $A=s\mathcal{I}+B$ where $s\in\mathbb{R}$ and $B\in\mathbb{R}^{[m,n]}$ is a nonnegative tensor.
@@ II-B $\mathcal{M}$-tensors and $\mathcal{H}^{+}$-tensors
$m_{i_{1}i_{2}\ldots i_{m}}=\left\{\begin{array}[]{cl}\left|A_{i_{1}i_{2}\ldots i_{m}}\right|&\left(i_{1}i_{2}\ldots i_{m}\right)=\left(i_{1}i_{1}\ldots i_{1}\right),\\ -\left|A_{i_{1}i_{2}\ldots i_{m}}\right|&\left(i_{1}i_{2}\ldots i_{m}\right)\neq\left(i_{1}i_{1}\ldots i_{1}\right).\end{array}\right.$
Notice that a tensor $A\in\mathbb{R}^{[m,n]}$ is an $\mathcal{H}$-tensor if its comparison tensor is an $\mathcal{M}$-tensor; it is a nonsingular $\mathcal{H}$-tensor, if its comparison tensor is a nonsingular $\mathcal{M}$-tensor. A nonsingular $\mathcal{H}$-tensor $A$ with all its diagonal elements $a_{ii\ldots i}>0$ is called an $\mathcal{H}^{+}$-tensor.
A tensor ${A}\in\mathbb{R}^{[m,n]}$ is called diagonally dominant if
@@
II-C $\mathcal{S}$-tensors and their properties
Here, we briefly recall the concept of an $\mathcal{S}$-tensor and we leverage their properties [16] to achieve some further results regarding non-homogeneous equation systems related to $\mathcal{S}$-tensors.
A tensor $A\in\mathbb{R}^{[m,n]}$ is an $\mathcal{S}$-tensor, if there exists a positive vector $x$ such that $Ax^{k-1}>\mathbf{0}$. Clearly, an $\mathcal{H}^{+}$-tensor is an $\mathcal{S}$-tensor.
@@ II-C $\mathcal{S}$-tensors and their properties
Notice indeed, that if $A$ is an $\mathcal{H}^{+}$-tensor or $\mathcal{M}$-tensor, then both conditions in Lemma 2 hold automatically, see [15, section 3]. Then, we have the following.
@@
If $A\in\mathbb{R}^{[m,n]}$ is an $\mathcal{H}^{+}$-tensor, then for every positive vector $b\mathbb{R}^{n}$ the tensor equation $Ax^{m-1}=b$ has a unique positive solution.
@@
where $b$ is a positive vector. We will show later that this equation system is relevant to the higher-order Lotka-Volterra model.
We will use the following Lemma to provide the result.
@@
Let $\mathbb{P}$ be a regular cone in an ordered Banach space $\mathbb{E}$ and $[u,v]\subset\mathbb{E}$ be a bounded ordererd interval. Suppose that $T:[u,v]\mapsto\mathbb{E}$ is an increasing continuous map which satisfies
@@
$x_{k+1}=T\left(x_{k}\right),\quad k=0,1,2,\ldots$
@@
Proof:
Let $A_{i}\in\mathbb{R}^{[i,n]}$. Following [16], for any tensor $A$, it can be formulated as $A_{i}=\alpha_{i}\mathcal{I}-\mathcal{B}_{i}+\mathcal{N}_{i}$, where $\alpha_{i}=\max\left\{0,(A_{i})_{1\cdots 1},(A_{i})_{2\cdots 2},\ldots,(A_{i})_{n\cdots n}\right\},\mathcal{B}_{i}$ is a nonnegative tensor whose diagonal entries $(\mathcal{B}_{i})_{jj\cdots j}=\alpha-$ $(A_{i})_{jj\ldots j}(j=1,2,\ldots,n)$ and nondiagonal entries equal to the opposite of the negative elements of $A_{i}$ other than diagonal elements, and $\mathcal{N}_{i}$ is a nonnegative tensor which includes the positive elements of $A_{i}$ other than diagonal elements.
Now, let $f(x)=\sum_{i=2}^{k}(\alpha_{i}\mathcal{I}+\mathcal{N}_{i})x^{i-1}$. We further define $g(x)=\sum_{i=2}^{k}\mathcal{B}_{i}x^{i-1}+b$. Now, since $b$ is a positive vector and $\mathcal{B}_{i}$ is nonnegative, we have that $g(x)$ is an increasing map on ++$\mathbb{R}^{n}_{++}$. The map $f(x):\mathbb{R}^{n}_{++}\mapsto\mathbb{R}^{n}_{++}$ is an increasing continuous map and thus it has an increasing continuous inverse $f^{-1}(x)$ on ++$\mathbb{R}^{n}_{++}$. Hence, a solution of the tensor equation (2) can be written as:
@@
Here, we choose a $t$ such that $\underline{\gamma}=1$. Let $\underline{x}=tv$. We notice that $t\leq w$ implies that $\underline{x}\leq\bar{x}$.
Similarly,
@@
This leads to $\tau{y_{j}}<[T(\tau y)]_{j}\leqslant[T(x)]_{j}=x_{j}=\tau{y_{j}}$, which is a contradiction. Thus, it must hold $\tau\geq 1$, which leads to $x\geq y$.
Now, let $\xi=\max\frac{x_{i}}{y_{i}}$. Then, $x\leqslant\xi y$ and $x_{j}=\xi y_{j}$ for some $j$. If $\xi>1$, we have
@@
Then, $\xi{y_{j}}>[T(\xi y)]_{j}\geq[T(x)]_{j}=x_{j}=\xi{y_{j}}$. This is a contradiction. Thus, it must hold $\xi\leq 1$, which leads to $y\geq x$. Combined with the above argument $x\geq y$, we have $x=y$ and thus the fixed point must be unique. ∎
Remark 1
The Theorem requires $b=\mathbf{1}$. However, the result is still very general. Consider the equation system (2) with an arbitrary positive $b$ and any arbitrary tensors $A_{i}$. Let $(\tilde{A}_{i})_{j_{1},\cdots,j_{i}}=\frac{(A_{i})_{j_{1},\cdots,j_{i}}}{b_{j_{1}}}$. The original equation system (2) is equivalent to $\tilde{A}_{k-1}x^{k-1}+\tilde{A}_{k-1}x^{k-2}+\cdots\tilde{A}_{2}x=\mathbf{1}$. Furthermore, for the proof, the way we construct $f(x),g(x)$ is different from the proof of Lemma 2 . Our approach can deal with a more complicated non-homogeneous case and doesn’t rely on condition ii) in Lemma 2. As a special case of theorem 1, the following Corollary further generalizes the result of Lemma 2.
Corollary 1
The homogeneous tensor equation $Ax^{k-1}=b$ with $b=\mathbf{1}$ has a unique positive solution if $A$ is an $\mathcal{S}$-tensor.
@@
This leads to $T(\bar{x})\leq\bar{x}$.
Next, we further look at
@@
Since $b$ is positive, $y$ must also be positive. This further implies that $T(\mathbf{0})=f^{-1}(b)=y\geq\mathbf{0}$. From Lemma 4, there must be at least one positive solution. We know from Lemma 4 that the iteration must converge to the boundary solution from below and if $x_{*}=x^{*}$, then the positive solution is unique. ∎
Remark 2
@@
II-D Polynomial complementarity problem
In tensor algebra, the polynomial complementarity problem is strongly related to the properties of tensors [18].
The polynomial complementarity problem (PCP) is defined by
$\begin{split}x&\geq\mathbf{0},\\ F(x)+q=\sum_{k=2}^{m}{A}_{k}x^{k-1}+q&\geq\mathbf{0},\\ x^{\top}(F(x)+q)=x^{\top}\left(\sum_{k=2}^{m}{A}_{k}x^{k-1}+q\right)&=\mathbf{0},\end{split}$ (12)

### v2

@@ Abstract
It is known that the effect of species’ density on species’ growth is non-additive in real ecological systems. This challenges the conventional Lotka-Volterra model, where the interactions are always pairwise and their effects are additive. To address this challenge, we introduce HOIs (Higher-Order Interactions) which are able to capture, for example, the indirect effect of one species on a second one correlating to a third species. Towards this end, we study a general higher-order Lotka-Volterra model. We provide an existence result of a positive equilibrium for a non-homogeneous polynomial equation system with the help of $\mathcal{S}$-tensors. Afterward, by utilizing the latter result, as well as the theory of monotone systems and results from the polynomial complementarity problem, we provide comprehensive results regarding the existence, uniqueness, and stability of the corresponding equilibrium. These results can be regarded as natural extensions of many analogous ones for the classical Lotka-Volterra model, especially in the case of full cooperation, competition among two factions, and pure competition. Finally, illustrative numerical examples are provided to highlight our contributions.
{IEEEkeywords}
@@ 1 Introduction
Since the Lotka-Volterra model is a polynomial system, computing a positive equilibrium is equivalent to solving a system of polynomial equations. With the development of the tensor algebra, a polynomial equation system can be efficiently captured by the tensor-vector product [14, 15, 16]. Researchers have mainly focused on homogeneous polynomial systems and have further shown that such systems may possess a unique positive solution for some particular classes of tensors and under some appropriate conditions. For example, [14] provides a result of the existence and uniqueness of a positive solution for $\mathcal{M}$-tensors, while [15] extends the results to $\mathcal{H}^{+}$-tensors, and [16] to $\mathcal{S}$-tensors, see section 2. However, most of these results are restricted to a homogeneous polynomial case and are not directly applicable to the Lotka-Volterra model.
@@
Notation: Throughout this paper, $\mathbb{R}$ denotes the set of real numbers, $\mathbb{R}_{+}$ is the set of nonnegative real numbers, and $\mathbb{R}_{++}$ the set of positive real numbers. For the similar notation $\mathbb{R}^{n}$ (+$\mathbb{R}^{n}_{+}$ or ++$\mathbb{R}^{n}_{++}$), the superscript stands for the dimension of the space. Whenever $a,b\in\mathbb{R}^{n}$, we use the notation $a\geq(\leq)b$ to denote that $a_{i}\geq(\leq)b_{i}$, for all $i=1,\ldots,n$; $a>(<)b$ to denote that $a_{i}>(<)b_{i}$, for all $i=1,\ldots,n$. This element-wise comparison is also used for matrices and tensors. For simplicity, the equilibrium point denotes both the equilibrium point itself and the vector constructed from the equilibrium point. For a vector $X=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}$, the vector’s $L_{\infty}$ norm is defined as $\|X\|_{\infty}=\max_{i}\left|x_{i}\right|$.
2 Tensor-based approach
In this section, we briefly introduce the concepts of tensors and polynomial complementarity problems that are useful later in the paper. We also propose some novel results, which are extensions of some known results in tensor algebra [14, 15, 16]. We will then show that all these results are very useful tools to analyze the higher-order Lotka-Volterra system.
2.1 Tensor
A tensor $T\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{k}}$ is a multidimensional array, where the order is the number of its dimensions $k$ and each dimension $n_{i}$, $i=1,\cdots,k$ is a mode of the tensor. A tensor is cubical if every mode has the same size, that is $T\in\mathbb{R}^{n\times n\times\cdots\times n}$. From now on we denote a $k$-th order $n$ dimensional cubical tensor as $T\in\mathbb{R}^{n\times n\times\cdots\times n}=\mathbb{R}^{[k,n]}$. Throughout this paper, a ‘tensor’ always refers to a cubical tensor. A tensor $T$ is called supersymmetric if $T_{j_{1}j_{2}\ldots j_{k}}$ is invariant under any permutation of the indices. The identity tensor $\mathcal{I}=\left(\delta_{i_{1}\cdots i_{m}}\right)$ is defined by
$\delta_{i_{1}\cdots i_{m}}=\begin{cases}1&\text{ if }i_{1}=i_{2}=\cdots=i_{m},\\ 0&\text{ otherwise.}\end{cases}$
@@ 2.1 Tensor
2.2 $\mathcal{M}$-tensors and $\mathcal{H}^{+}$-tensors
Here, we briefly recall the concepts of $\mathcal{M}$-tensor and $\mathcal{H}^{+}$-tensor introduced in [15, 26, 27, 28]. A tensor $A\in\mathbb{R}^{[m,n]}$ is an $\mathcal{M}$-tensor [26, 27] if it can be represented as $A=s\mathcal{I}-B$, where $\mathcal{I}\in\mathbb{R}^{[m,n]}$ is the identity tensor, $B\in\mathbb{R}^{[m,n]}$ is a nonnegative tensor (i.e., each entry of $B$ is nonnegative), and $s\geq\rho(B)$. Furthermore, $A$ is called a nonsingular $\mathcal{M}$-tensor if $s>\rho(B)$. A tensor $A\in\mathbb{R}^{[m,n]}$ is a Metzler tensor [28] if it can be represented as $A=s\mathcal{I}+B$ where $s\in\mathbb{R}$ and $B\in\mathbb{R}^{[m,n]}$ is a nonnegative tensor. In other words, a Metzler tensor is a tensor whose off-diagonal elements are all nonnegative.
@@ 2.2 $\mathcal{M}$-tensors and $\mathcal{H}^{+}$-tensors
$m_{i_{1}i_{2}\ldots i_{m}}=\left\{\begin{array}[]{cl}\left|A_{i_{1}i_{2}\ldots i_{m}}\right|,&\left(i_{1}i_{2}\ldots i_{m}\right)=\left(i_{1}i_{1}\ldots i_{1}\right),\\ -\left|A_{i_{1}i_{2}\ldots i_{m}}\right|,&\left(i_{1}i_{2}\ldots i_{m}\right)\neq\left(i_{1}i_{1}\ldots i_{1}\right).\end{array}\right.$
Notice that a tensor $A\in\mathbb{R}^{[m,n]}$ is an $\mathcal{H}$-tensor [15] if its comparison tensor is an $\mathcal{M}$-tensor; it is a nonsingular $\mathcal{H}$-tensor, if its comparison tensor is a nonsingular $\mathcal{M}$-tensor. A nonsingular $\mathcal{H}$-tensor $A$ with all its diagonal elements $a_{ii\ldots i}>0$ is called an $\mathcal{H}^{+}$-tensor [15].
A tensor ${A}\in\mathbb{R}^{[m,n]}$ is called diagonally dominant [29] if
@@
2.3 $\mathcal{S}$-tensors and their properties
Here, we briefly recall the concept of an $\mathcal{S}$-tensor and we leverage their properties [16] to achieve some further results regarding non-homogeneous systems related to $\mathcal{S}$-tensors.
A tensor $A\in\mathbb{R}^{[m,n]}$ is an $\mathcal{S}$-tensor [16], if there exists a positive vector $x$ such that $Ax^{k-1}>\mathbf{0}$. Clearly, an $\mathcal{H}^{+}$-tensor is an $\mathcal{S}$-tensor.
@@ 2.3 $\mathcal{S}$-tensors and their properties
Notice indeed, that if $A$ is an $\mathcal{H}^{+}$-tensor or an $\mathcal{M}$-tensor, then both conditions in Lemma 2 hold automatically, see [15, section 3]. Then, we have the following.
@@
If $A\in\mathbb{R}^{[m,n]}$ is an $\mathcal{H}^{+}$-tensor, then for every positive vector $b\in\mathbb{R}_{++}^{n}$, the tensor equation $Ax^{m-1}=b$ has a unique positive solution.
@@
where $b$ is a positive vector. We will show later that this equation system is relevant to the higher-order Lotka-Volterra model. The following Lemma will thus be important.
@@
Let $\mathbb{P}$ be a regular cone in an ordered Banach space $\mathbb{E}$ and $[u,v]\subset\mathbb{E}$ be a bounded ordererd interval. Suppose that $T:[u,v]\rightarrow\mathbb{E}$ is an increasing continuous map which satisfies
@@
$x_{k+1}=T\left(x_{k}\right),\quad k=0,1,2,\ldots,$
@@
We now show that, under certain conditions, the non-homogeneous tensor equation (2) has a unique solution.
@@
Proof 2.1.
Let $A_{i}\in\mathbb{R}^{[i,n]}$. Following [16], any tensor $A_{i}$ can be rewritten as $A_{i}=\alpha_{i}\mathcal{I}-\mathcal{B}_{i}+\mathcal{N}_{i}$, where $\alpha_{i}=\max\left\{0,(A_{i})_{1\cdots 1},(A_{i})_{2\cdots 2},\ldots,(A_{i})_{n\cdots n}\right\},\mathcal{B}_{i}$ is a nonnegative tensor whose diagonal entries $(\mathcal{B}_{i})_{jj\cdots j}=\alpha-$ $(A_{i})_{jj\ldots j}\;(j=1,2,\ldots,n)$ and nondiagonal entries equal to the opposite of the negative elements of $A_{i}$ other than diagonal elements, and $\mathcal{N}_{i}$ is a nonnegative tensor which includes the positive elements of $A_{i}$ other than diagonal elements.
Next, let $f(x)=\sum_{i=2}^{k}(\alpha_{i}\mathcal{I}+\mathcal{N}_{i})x^{i-1}$. We further define $g(x)=\sum_{i=2}^{k}\mathcal{B}_{i}x^{i-1}+b$. Now, since $b$ is a positive vector and $\mathcal{B}_{i}$ is nonnegative, we have that $g(x)$ is an increasing map on ++$\mathbb{R}^{n}_{++}$. The map $f(x):\mathbb{R}^{n}_{++}\rightarrow\mathbb{R}^{n}_{++}$ is an increasing continuous map and thus it has an increasing continuous inverse $f^{-1}(x)$ on ++$\mathbb{R}^{n}_{++}$. Hence, a solution of the tensor equation (2) can be written as:
@@
Here, we choose a $t$ such that $\underline{\gamma}=1$. Letting $\underline{x}=tv$ we notice that $t\leq w$ implies that $\underline{x}\leq\bar{x}$. Analogously,
@@
This leads to $\tau{y_{j}}<[T(\tau y)]_{j}\leqslant[T(x)]_{j}=x_{j}=\tau{y_{j}}$, which is a contradiction. Thus, it must hold $\tau\geq 1$, which leads to $x\geq y$. So, let $\xi=\max\frac{x_{i}}{y_{i}}$. Then, $x\leqslant\xi y$ and $x_{j}=\xi y_{j}$ for some $j$. If $\xi>1$, we have
@@
Then, $\xi{y_{j}}>[T(\xi y)]_{j}\geq[T(x)]_{j}=x_{j}=\xi{y_{j}}$. This is again a contradiction. Thus, it must hold $\xi\leq 1$, which leads to $y\geq x$. Combined with the above argument $x\geq y$, we have $x=y$. Therefore the fixed point is unique.
Remark 2.2.
Theorem 1 requires $b=\mathbf{1}$. However, the result is still very general. Consider (2) with an arbitrary positive $b$ and any arbitrary tensors $A_{i}$. Let $(\tilde{A}_{i})_{j_{1},\cdots,j_{i}}=\frac{(A_{i})_{j_{1},\cdots,j_{i}}}{b_{j_{1}}}$. Then, (2) is equivalent to $\tilde{A}_{k-1}x^{k-1}+\tilde{A}_{k-1}x^{k-2}+\cdots\tilde{A}_{2}x=\mathbf{1}$. Furthermore, the way we construct $f(x),g(x)$ for the proof of Theorem 1 is different from the proof of Lemma 2. Our approach can deal with a more complicated non-homogeneous case and doesn’t rely on condition ii) in Lemma 2. As a special case of Theorem 1, the following Corollary further generalizes the result of Lemma 2.
Corollary 2.
If $A$ is an $\mathcal{S}$-tensor, then the homogeneous tensor equation $Ax^{k-1}=b$ with $b=\mathbf{1}$ has a unique positive solution.
@@
This leads to $T(\bar{x})\leq\bar{x}$. Next, we further look at
@@
Since $b$ is positive, $y$ must also be positive. This further implies that $T(\mathbf{0})=f^{-1}(b)=y\geq\mathbf{0}$. From Lemma 4, there must be at least one positive solution. We know from Lemma 4 that the iteration must converge to the boundary solution from below and if $x_{*}=x^{*}$, then the positive solution is unique.
Remark 2.4.
@@
2.4 Polynomial complementarity problem
In tensor algebra, the polynomial complementarity problem is strongly related to the properties of tensors [18]. The polynomial complementarity problem (PCP) is defined by
$\begin{split}x&\geq\mathbf{0},\\ F(x)+q=\sum_{k=2}^{m}{A}_{k}x^{k-1}+q&\geq\mathbf{0},\\ x^{\top}(F(x)+q)=x^{\top}\left(\sum_{k=2}^{m}{A}_{k}x^{k-1}+q\right)&=0,\end{split}$ (12)